# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ Appendix I Noise amplitude in the transformed reward estimate
SFT $\rightarrow$ offpolicy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1874}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0000}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1904}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1871}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1903}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ 0.01 1e-6 160
SFT $\rightarrow$ offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1872}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1907}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1900}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1922}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 0.01 1e-6 480
@@
offpolicy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1830}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1858}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1714}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1755}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ 2.5 3e-7 160
offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1822}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1857}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1735}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0037}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1740}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0008}}$ 10 3e-7 320
@@
SFT $\rightarrow$ offpolicy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1918}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1941}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1914}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1936}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 10 3e-7 320
SFT $\rightarrow$ offpolicy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1912}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1939}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1910}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1938}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ 2.5 3e-7 764

### v2

@@ Appendix I Noise amplitude in the partition function estimate for the initial reward
SFT $\rightarrow$ off-policy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1874}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0000}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1904}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1871}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0006}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1903}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ 0.01 1e-6 160
SFT $\rightarrow$ off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1872}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1907}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1900}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1922}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 0.01 1e-6 480
@@
off-policy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1830}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1858}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1714}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1755}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0007}}$ 2.5 3e-7 160
off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1822}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1857}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0004}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1735}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0037}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1740}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0008}}$ 10 3e-7 320
@@
SFT $\rightarrow$ off-policy–best ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1918}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1941}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1914}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1936}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0001}}$ 10 3e-7 320
SFT $\rightarrow$ off-policy–random ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1912}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0002}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1939}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1910}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0003}}$ / ${\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.1938}\,\text{\tiny$\pm$}\,\text{\tiny{\color[rgb]{0.2,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.2,0.2,0.2}\pgfsys@color@gray@stroke{0.2}\pgfsys@color@gray@fill{0.2}0.0005}}$ 2.5 3e-7 764